OpenStudy (anonymous):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10−6 cm/s B. 8.889 × 10−5 cm/s C. 1.152 × 10−5 cm/s D. 1.152 × 10−7 cm/s

OpenStudy (bibby):

"3.2 × 10−4 " using the following latex "\large 3.2 × 10^{−4} we get: "\(\large 3.2 × 10^{−4} \)

OpenStudy (bibby):

oh god what

OpenStudy (anonymous):

omg yeah ikr it's a mess

OpenStudy (bibby):

anyhow convert miles per hour to miles per second, then we can convert miles to meters to centimeters

OpenStudy (anonymous):

problem is , i don't know how .. ugh more math to study cries

OpenStudy (bibby):

I have a used tissue somewhere here if you need it

OpenStudy (anonymous):

oh my thanks yes , I'm running out >_<

OpenStudy (bibby):

ok. here's an example of dimensional analysis. if we wanted to convert 320km to miles: we know that 1 mile is 1.60934km thanks to google. Next we multiply by a proportion that is equal to 1 that allows us to "cancel" the units. \[\huge \frac{320\cancel{km}}{}*\frac{1mi}{1.60934\cancel{km}}=...\]

OpenStudy (anonymous):

I've never been so confused in my life

OpenStudy (bibby):

Google/youtube dimensional analysis. Basically if we want to convert 30 minutes to hours we would do the following \(\large \frac{30 \cancel{min}}{1} * \frac{1 hour}{60 \cancel{min}}=\frac{30}{60}=0.5hours\)

OpenStudy (kirbykirby):

I already posted this in your previous post http://openstudy.com/users/kirbykirby#/updates/540a0470e4b0d8ead1bcb9e8

OpenStudy (kirbykirby):

@bibby Hmm I thought the \(m\) was for metres. @kewlkatcutie Do you know if they are using metres or miles?

OpenStudy (bibby):

the m is for meters. I gave an irrelevant example to demonstrate dimensional analysis @kirbykirby

OpenStudy (anonymous):

It doesn't say ; _ ;

OpenStudy (kirbykirby):

Oh ok I see :)

OpenStudy (anonymous):

Is the correct answer D ?

OpenStudy (kirbykirby):

What did you do to find that? All you need to do is convert the metres into centimetres, and then the hours into seconds. 1 m = 100 cm 1 hour = 60 min..... but 1 min = 60 seconds. So 1 hour = (60)(60) = 3600 seconds. I posted the right way though to do the dimensional analysis that shows that the units are correct

OpenStudy (kirbykirby):

Follow @bibby's examples up there.. it's maybe more clear from them what the idea behind dimensional analysis is

OpenStudy (anonymous):

so .. \[\frac{ 100 }{ 3600 }\] .. ?

OpenStudy (bibby):

\[\huge \frac{3.2*10^{-4}m}{1\cancel{hr}}*\frac{1\cancel{hr}}{60 \cancel{\min}}*\frac{1\cancel{\min}}{60 \sec}\] that's how you get meters per second